Displaying report 1-1 of 1.
Reports until 16:00, Thursday 03 April 2025
GQuEST Output Filter Cavities
Torrey Cullen - posted 16:00, Thursday 03 April 2025 (12224)
Stable Piezo Lock in OFC3

I set up the piezo control for the third filter cavity. This is to allow a way to scan cavity 3 while keeping OFC2 laser locked on 1550. The scan works and after tuning the controller a little it provide one of the best piezo locks we've seen in these cavities. Note that this new cavity is with the new mirror mounts and the nac2125 from CTS piezo. More analysis is needed but this piezo seems promising. For now I want to align the OFC3 1550 path, which I now should have all the tools for. Will come back to this new piezo analysis later.

Displaying report 1-1 of 1.